### v2

# Geometric Time-Dependent Density Functional Theory

Abstract
We provide a new formulation of Time-Dependent Density Functional Theory (TDDFT) based on the geometric structure of the set of states constrained to have a fixed density. Orbital-free TDDFT is formulated using a hydrodynamics equation involving a new density-to-current functional map. In the corresponding Kohn–Sham equation, the density is reproduced using a non-local operator. Finally, we present numerical simulations for one-dimensional soft-Coulomb systems.
Density Functional Theory (DFT) is very successful in describing electrons at equilibrium, that is, close to their ground states. It is nowadays the reference computational method for large molecules and solids, both in academic research and industrial applications. By contrast, Time-Dependent Density Functional Theory (TDDFT) [1, 2, 3] is less developed and still suffers from important limitations [4, 5, 6]. It is however the only viable computational approach for complex systems at present, and the demand for novel concepts and approximations in TDDFT is therefore considerable. Practical applications of TDDFT include charge transfer in complex molecules [7], electronic response of large systems [8], and ultrafast phenomena in attosecond physics [9]. In this Letter we open a new avenue in TDDFT. Based on geometric concepts, we introduce a new exact scheme that seems to better describe systems driven out of equilibrium. We also present 1D numerical simulations that illustrate the potential power of our new approach, when used as a correction to adiabatic density functionals.
Geometric orbital-free theory.
We consider the usual $N$-electron Hamiltonian
$\widehat{H}(t):=\widehat{T}+\widehat{V}_{\rm ee}+\widehat{V}_{\rm ext}(t)$ (1)
where $\widehat{T}:=-\sum_{j=1}^{N}\nabla^{2}_{\mathbf{r}_{j}}/2$ is the kinetic-energy operator, $\widehat{V}_{\rm ee}$ is the electron-electron interaction and $\widehat{V}_{\rm ext}(t):=\sum_{j=1}^{N}V_{\rm ext}(t,\mathbf{r}_{j})$ is a time-dependent external potential. The reference $N$-particle Schrödinger equation is
$i\partial_{t}\Psi^{\rm S}(t)=\widehat{H}(t)\,\Psi^{\rm S}(t),\quad\Psi^{\rm S}(0)=\Psi_{0}.$ (2)
In this first section, we focus on orbital-free TDDFT, where the high-dimensional linear equation (2) is replaced by an equivalent nonlinear equation involving only the one-particle density
$\rho_{\Psi}(\mathbf{r}):=\bigg\langle\Psi,\sum_{j=1}^{N}\delta(\mathbf{r}-\mathbf{r}_{j})\Psi\bigg\rangle.$ (3)
To this end, we consider modified Schrödinger equations of the form
$i\partial_{t}\Psi(t)=\Big(\widehat{H}(t)+\widehat{F}[\Psi_{0},V_{\rm ext},\rho](t)\Big)\Psi(t),\quad\Psi(0)=\Psi_{0},$ (4)
where the correction term $\widehat{F}[\Psi_{0},V_{\rm ext},\rho](t)$ is such that the density $\rho_{\Psi(t)}(\mathbf{r})$ of the solution is equal to a density $\rho(t,\mathbf{r})$ that is prescribed in advance. The correction term is a functional of the density $\rho$, of the time-dependent external potential $V_{\rm ext}$ and of the initial state $\Psi_{0}$ which is of course assumed to have density $\rho(0,\mathbf{r})$. To ensure causality, $\widehat{F}[\Psi_{0},V_{\rm ext},\rho](t)$ is assumed to only depend on $\rho(s,\mathbf{r})$ and $V_{\rm ext}(s,\mathbf{r})$ for past times $0\leqslant s\leqslant t$.
There are many possible choices for $\widehat{F}$. Based on geometric considerations, we will construct in this section a functional for which the density $\rho^{\rm S}:=\rho_{\Psi^{\rm S}}$ of the Schrödinger wavefunction $\Psi^{\rm S}$ in (2) is the only solution to the implicit equation $\widehat{F}[\Psi_{0},V_{\rm ext},\rho]=0$. This is how (2) can be reformulated in terms of the density only. In the next section, we consider the corresponding Kohn–Sham theory which provides $\rho^{\rm S}$ in terms of orbitals obtained from a non-interacting Hamiltonian with a similar but different correction $\widehat{F}$ that is non-zero for $\rho=\rho^{\rm S}$.
Standard TDDFT can be reinterpreted as the choice $\widehat{F}^{\rm VP}[\Psi_{0},V_{\rm ext},\rho](t):=\sum_{j=1}^{N}V(t,\mathbf{r}_{j})$ with a potential functional $V=V[\Psi_{0},V_{\rm ext},\rho]$. This can be achieved by requiring the action to be stationary with respect to wavefunction variations at fixed density [10, Sec. III.A], which we refer to as the Variational Principle (VP). The Runge–Gross theorem [1] states that $V$ is unique modulo a time-dependent constant, hence $\rho^{\rm S}$ is the unique solution to $V[\Psi_{0},V_{\rm ext},\rho]=0$. Since $V[\Psi_{0},V_{\rm ext},\rho]=V[\Psi_{0},0,\rho]-V_{\rm ext}$, this last condition can be formulated with the universal functional $V[\Psi_{0},\rho]:=V[\Psi_{0},0,\rho]$ as
$V[\Psi_{0},\rho](t,\mathbf{r})=V_{\rm ext}(t,\mathbf{r})+C(t)\quad\Longleftrightarrow\quad\rho=\rho^{\rm S},$ (5)
where $C(t)$ is a time-dependent constant.
Figure 1: In the Geometric Principle, the velocity $\partial_{t}\Psi(t)$ is the projection of $i\widehat{H}(t)\Psi(t)$ onto the tangent space $\mathcal{T}_{\Psi(t)}\mathcal{M}_{\rho}$. The normal space is given by (6) and one arrives at Eq. (8).
In our new Geometric Principle (GP) we do not demand the action to be stationary, but instead require that the norm of the correction, $\|\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]\Psi(t)\|=\|i\partial_{t}\Psi(t)-\widehat{H}(t)\Psi(t)\|$, be the smallest it can be at all times. For simplicity, we first consider a prescribed time-independent density $\rho(\mathbf{r})$. The wavefunction $\Psi(t)$ in (4) then evolves on the set $\mathcal{M}_{\rho}$ of all wavefunctions having this prescribed $\rho$ and the velocity $\partial_{t}\Psi(t)$ must belong to the tangent space $\mathcal{T}_{\Psi(t)}\mathcal{M}_{\rho}$ at all times. As shown in Fig. 1, the correction term has minimal norm when $\partial_{t}\Psi(t)$ is the projection of $i\widehat{H}(t)\Psi(t)$ onto the tangent space, that is, when $i\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]\Psi(t)$ is in the normal space $\mathcal{N}_{\Psi(t)}:=(\mathcal{T}_{\Psi(t)}\mathcal{M}_{\rho})^{\perp}$. By simple calculations [10, Sec. II.B], the normal space is
$\mathcal{N}_{\Psi}=\bigg\{\sum_{j=1}^{N}w(\mathbf{r}_{j})\Psi,\ w:\mathbb{R}^{3}\to\mathbb{R}\bigg\}$ (6)
and the new correction term is thus
$\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]\Psi(t):=i\sum_{j=1}^{N}W(t,\mathbf{r}_{j})\Psi(t),$ (7)
with a real-valued function $W(t,\mathbf{r})$. For a prescribed time-dependent density, the displacement of the tangent space in the normal direction only depends on $\partial_{t}\rho(t)$ and does not change the form of the correction in (7[10, Sec. II.C]. By construction, $W(t,\mathbf{r})$ only depends on $\Psi(t)$, $V_{\rm ext}(t)$ and $\partial_{t}\rho(t)$ and is thus causal. Hence the modified Schrödinger equation (4) is
$i\partial_{t}\Psi(t)=\left(\widehat{H}(t)+i\widehat{W}(t)\right)\Psi(t),$ (8)
where $\widehat{W}(t):=\sum_{j=1}^{N}W(t,\mathbf{r}_{j})$ and $W=W[\Psi_{0},V_{\rm ext},\rho]$ is a real-valued functional.
Equation (8) looks like a Schrödinger equation with a complex potential, similar to those considered in TDDFT for describing resonances [11, 12, 13, 14, 15] or open systems [16, 17, 18]. However, here, $W$ is chosen so that $\rho_{\Psi(t)}=\rho(t)$ and thus the resulting $\Psi(t)$ stays normalized for all times. This means that $\langle\Psi(t),\widehat{W}(t)\Psi(t)\rangle=0$, which implies that Eq. (8) can be rewritten in commutator form
$i\partial_{t}\Psi(t)=\bigg(\widehat{H}(t)+i\Big[\widehat{W}(t)\,,\,|\Psi(t)\rangle\langle\Psi(t)|\Big]\bigg)\Psi(t).$ (9)
For any fixed $\widehat{W}(t)$ and $\Psi(t)$, the modified Hamiltonian in (9) is Hermitian, and thus manifestly norm preserving.
The idea of constraining a Schrödinger dynamics using a projection on the tangent space, often attributed to McLachlan [19], can be traced back to Dirac [20] and Frenkel [21]. Depending on the considered constraints, the GP can lead to the same solution as the VP, or to completely different solutions [22, 23, 24, 25, 26, 27]. For a fixed density the GP and VP corrections and wavefunctions are very different [10, Secs. IV.A, VIII.C-D].
In [10, Thm. 3] we give a rigorous proof of the existence and uniqueness of $W[\Psi_{0},V_{\rm ext},\rho]$ in finite dimension. In Appendix A we rigorously prove a Runge–Gross theorem for Eq. (8) with smooth potentials. Hence $W[\Psi_{0},V_{\rm ext},\rho]$ is uniquely defined, when it exists, and $\rho^{\rm S}$ is the unique solution to $\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]=0$, that is,
$W[\Psi_{0},V_{\rm ext},\rho]=0\quad\Longleftrightarrow\quad\rho=\rho^{\rm S}.$ (10)
The function $W$ is very different in nature from the potential $V$ used in standard TDDFT. In particular, the potential $V$ often contains high peaks and steps to force the particles to move to the desired region of space [28]. This only works thanks to a complicated interplay between $V$ and the kinetic-energy operator $\widehat{T}$. Indeed, if we remove $\widehat{T}$, then adding a real-valued local potential $V$ modifies the phase of the wavefunction but not the density. In constrast, the function $W$ appears directly in the continuity equation associated with (8) or (9),
$\partial_{t}\rho_{\Psi(t)}(\mathbf{r})+\nabla\cdot\mathbf{j}_{\Psi(t)}(\mathbf{r})=2\big(\mathcal{L}_{\Psi(t)}W(t)\big)(\mathbf{r})$ (11)
with the linear operator
$\big(\mathcal{L}_{\Psi}w\big)(\mathbf{r})=\rho_{\Psi}(\mathbf{r})w(\mathbf{r})+\int_{\mathbb{R}^{3}}\rho^{(2)}_{\Psi}(\mathbf{r},\mathbf{r}^{\prime})w(\mathbf{r}^{\prime})\,\mathrm{d}\mathbf{r}^{\prime},$ (12)
where $\mathbf{j}_{\Psi}$ and $\rho_{\Psi}^{(2)}$ are the current and the two-particle density of $\Psi$. The linear operator $\mathcal{L}_{\Psi}$ is expected to be invertible in all cases of physical interest [10, Sec. VII.B]. The interpretation of (11) is that $W$ gives rise to sources and sinks [12], which is a more direct and flexible way of changing the density. In fact, on a lattice, densities moving too fast are not $V$-representable due to the finite speed of propagation [13] but can perfectly be $W$-representable [10, Secs. VII.C and VII.D].
Let us now re-interpret Eq. (10) in light of the continuity equation (11). From the functional $W[\Psi_{0},V_{\rm ext},\rho]$ we obtain a unique wavefunction $\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t)$ by solving (8) or (9), as well as the corresponding geometric density-to-current functional
$\mathbf{j}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t,\mathbf{r}):=\mathbf{j}_{\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t)}(\mathbf{r}),$ (13)
the current of the wavefunction $\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t)$. It plays a central role in our new theory because, from the invertibility of $\mathcal{L}_{\Psi}$ in (11), we can rewrite Eq. (10) as
$\partial_{t}\rho(t,\mathbf{r})+\nabla\cdot\mathbf{j}^{\rm GP}\big[\Psi_{0},V_{\rm ext},\rho\big](t,\mathbf{r})=0\quad\Longleftrightarrow\quad\rho=\rho^{\rm S}.$ (14)
In other words, the sought-after $\rho^{\rm S}$ is the unique density that satisfies the standard continuity equation with the geometric current $\mathbf{j}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t,\mathbf{r})$. For $\rho=\rho^{\rm S}$, we recover the Schrödinger wavefunction and current, $\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho^{\rm S}]=\Psi^{\rm S}$ and $\mathbf{j}^{\rm GP}\big[\Psi_{0},V_{\rm ext},\rho^{\rm S}\big]=\mathbf{j}_{\Psi^{\rm S}(t)}$. Equation (14) is our new geometric formulation of orbital free TDDFT. It is analogous to the van Leeuwen equation [29] in standard TDDFT. Note that the latter is second-order in time whereas (14) is first-order.
Geometric Kohn–Sham theory.
In geometric Kohn–Sham theory, the exact density $\rho^{\rm S}$ is reproduced using $N$ non-interacting electrons described with a Slater determinant $\Phi(t):=(N!)^{-1/2}\det(\varphi_{n}(t,\mathbf{r}_{k}))$ with the orbitals $\varphi_{n}$ solving the equation
$i\partial_{t}\varphi_{n}(t,\mathbf{r})\!=\!\Big(-\frac{\nabla^{2}}{2}+V_{\rm ext}(t,\mathbf{r})+i\big[W^{\rm KS}(t),\gamma_{\Phi(t)}\big]\Big)\varphi_{n}(t,\mathbf{r}),$ (15)
where $\gamma_{\Phi(t)}(\mathbf{r},\mathbf{r}^{\prime}):=\sum_{n=1}^{N}\varphi_{n}(t,\mathbf{r})\varphi_{n}(t,\mathbf{r}^{\prime})^{\dagger}$ is the one-particle density matrix. The correction term is of commutator form as in (9) with $W^{\rm KS}$ a functional to be specified so that the exact Schrödinger density is reproduced, i.e. $\rho_{\Phi(t)}=\rho^{\rm S}(t)$. To obtain a predictive scheme, this functional should of course not depend on the unknown density $\rho^{\rm S}$.
For an initial Slater determinant $\Phi_{0}$, we can apply the theory developed before with $\widehat{V}_{\rm ee}=0$ and obtain the corresponding non-interacting functional $W_{0}[\Phi_{0},V_{\rm ext},\rho]$. If we pick $\rho=\rho^{\rm S}$, that is, solve (15) with $W^{\rm KS}=W_{0}[\Phi_{0},V_{\rm ext},\rho^{\rm S}]$, where $\Phi_{0}$ has the same density as the initial state $\Psi_{0}$, we obtain the trajectory of Slater determinants reproducing the exact Schrödinger density $\rho^{\rm S}$. In practice we do not know $\rho^{\rm S}$ and we can only conclude that the sought-after functional $W^{\rm KS}$ in Eq. (15) must coincide with $W_{0}[\Phi_{0},V_{\rm ext},\rho^{\rm S}]$ for (and only for) $\rho_{\Phi}=\rho^{\rm S}$. There are several choices for $W^{\rm KS}$.
Recall that in the Kohn–Sham scheme of standard TDDFT, the universal Hartree-exchange-correlation (Hxc) potential functional is defined as [31, Sec. 1.3]
$V_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]:=V_{0}[\Phi_{0},\rho]-V[\Psi_{0},\rho],$ (16)
where $V_{0}[\Phi_{0},\rho]$ is the universal Kohn–Sham potential and $V[\Psi_{0},\rho]$ the universal interacting potential functional as above. In the same spirit, we define here
$\widetilde{W}^{\rm KS}[\Psi_{0},\Phi_{0},V_{\rm ext},\rho]:=W_{0}[\Phi_{0},V_{\rm ext},\rho]-W[\Psi_{0},V_{\rm ext},\rho].$ (17)
Let us call $\Phi(t)$ the solution to the nonlinear equation (15) with $W^{\rm KS}=\widetilde{W}^{\rm KS}[\Psi_{0},\Phi_{0},V_{\rm ext},\rho_{\Phi}]$. By uniqueness in Thm. 1 (b) below, we know that this $W^{\rm KS}$ must be the one reproducing $\rho_{\Phi}$, which is by definition $W_{0}[\Phi_{0},V_{\rm ext},\rho_{\Phi}]$. But then from (17) we obtain $W[\Psi_{0},V_{\rm ext},\rho_{\Phi}]=0$ and therefore $\rho_{\Phi}=\rho^{\rm S}$ as we wanted.
The functional introduced in Eq. (17) is a natural extension of Eq. (16) but it implicitly depends on the external potential $V_{\rm ext}$. To turn it into a universal functional we can replace $V_{\rm ext}$ by $V[\Psi_{0},\rho]$ since for $\rho^{\rm S}$ we will have equality by (5). By definition any density $\rho$ is exactly reproduced with $V[\Psi_{0},\rho]$, hence no $W$ is necessary and we have $W[\Psi_{0},V[\Psi_{0},\rho],\rho]=0$. Thus, Eq. (17) leads to the universal geometric Kohn–Sham functional
$W^{\rm KS}[\Psi_{0},\Phi_{0},\rho]:=W_{0}\big[\Phi_{0},V[\Psi_{0},\rho],\rho\big],$ (18)
that is now independent of $V_{\rm ext}$. The solution $\Phi(t)$ to the nonlinear equation (15) with $W^{\rm KS}=W^{\rm KS}[\Psi_{0},\Phi_{0},\rho_{\Phi}]$ has the desired density $\rho^{\rm S}$. Indeed, by uniqueness in Thm. 1 (b) we know that this $W^{\rm KS}$ must be the one reproducing $\rho_{\Phi}$ in the external potential $V_{\rm ext}$, that is, $W_{0}\big[\Phi_{0},V[\Psi_{0},\rho_{\Phi}],\rho_{\Phi}\big]=W_{0}[\Phi_{0},V_{\rm ext},\rho_{\Phi}]$. But this equality means that the two potentials $V[\Psi_{0},\rho_{\Phi}]$ and $V_{\rm ext}$ yield the same $\rho_{\Phi}$ in presence of this $W^{\rm KS}$. By uniqueness of $V$ at given $W$ as stated in Thm. 1 (a) below, we conclude that $V[\Psi_{0},\rho_{\Phi}]=V_{\rm ext}+C(t)$. By (5) we finally obtain $\rho_{\Phi}=\rho^{\rm S}$ as desired.
We have introduced two nonlinear functionals to be used in the geometric Kohn–Sham scheme (15), that reproduce the exact Schrödinger density $\rho^{\rm S}$. There are other possibilities. The choice of a given functional should be motivated by how easily approximations can be made. This is not the focus of this Letter.
For systems close to equilibrium, standard adiabatic TDDFT approximations work well. However, this is not the case far from equilibrium. It is thus natural to use our new geometric theory to correct the limitations of the existing adiabatic approximations. For any given approximation $V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]$ to the exact Hxc potential in Eq. (16), we define the corresponding Kohn–Sham scheme with geometric correction as
$i\partial_{t}\varphi_{n}(t,\mathbf{r})=\Big(-\frac{\nabla^{2}}{2}+V_{\rm ext}(t,\mathbf{r})+V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho_{\Phi}](t,\mathbf{r})\\ +i\big[W^{\rm corr}(t)\,,\,\gamma_{\Phi(t)}\big]\Big)\varphi_{n}(t,\mathbf{r}).$ (19)
We can repeat everything we said before in the absence of $V^{\rm app}_{\rm Hxc}$. Namely, Eq. 19 will reproduce the exact Schrödinger density $\rho^{\rm S}$ if we take $W^{\rm corr}=\widetilde{W}^{\rm KS}$ with $V_{\rm ext}$ replaced by $V_{\rm ext}+V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]$ inside $W_{0}$ in the right-hand side of Eq. (17) or $W^{\rm corr}=W^{\rm KS}$ with $V[\Psi_{0},\rho]$ replaced by $V[\Psi_{0},\rho]+V^{\rm app}_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]$ in Eq. (18). If $V^{\rm app}_{\rm Hxc}$ is a good adiabatic approximation, we expect $W^{\rm corr}$ to be very small when close to a ground state and to only be significant in non-adiabatic situations. In the next paragraph we consider the case of two-electron spin-singlet states, for which Eq. (19) can be solved explicitly if $V_{\rm ext}+V^{\rm app}_{\rm Hxc}$ is the exact adiabatic Kohn–Sham potential for $\rho^{\rm S}$.
Two-electron spin-singlet.
[…]
$W$ $=\frac{\partial_{t}\rho^{\rm S}+\nabla\cdot(\rho^{\rm S}\nabla\theta)}{2\rho^{\rm S}},$ (20a)
$V$ $=-\partial_{t}\theta-\frac{|\nabla\theta|^{2}}{2}+\frac{\Delta\sqrt{\rho^{\rm S}}}{2\sqrt{\rho^{\rm S}}}.$ (20b)
In standard time-dependent Kohn–Sham theory where $W=0$, $\theta$ is obtained by solving (20a), and $V$ is given by (20b) with that $\theta$. For the geometric Kohn–Sham equation (15) where $V=V_{\rm ext}$ the situation is reversed: One should first solve (20b) to find $\theta$ and then $W$ is given by (20a). Note in passing that for the universal functional of Eq. (18) we take $V=V[\Psi_{0},\rho]$ and solve (20) with a general $\rho$ in place of $\rho^{\rm S}$. There are many solutions to (20). An important observation is that they can be parametrized by $\theta$, leading to explicit formulas for $V$ and $W$ from (20). The simplest case $\theta\equiv 0$ yields the solution
$V^{\text{eaKS}}=\frac{\Delta\sqrt{\rho^{\mathrm{S}}}}{2\sqrt{\rho^{\mathrm{S}}}},\qquad W^{\text{na}}=\frac{\partial_{t}\rho^{\mathrm{S}}}{2\rho^{\mathrm{S}}}.$ (21)
The potential $V^{\rm eaKS}$ is the exact adiabatic Kohn–Sham potential, defined as the total potential for which the time-dependent Kohn–Sham system has ground-state density $\rho^{\rm S}(t)$ for all $t$, and $W^{\text{na}}$ is the corresponding non-adiabatic geometric correction. We emphasize that we here apply the adiabatic approximation to the total exact Kohn–Sham potential $V^{\rm eKS}:=V_{0}[\Phi_{0},\rho^{\rm S}]$. It is not the usual adiabatic approximation, normally applied only to the Hxc potential in (16) [10, Sec. VIII.D]. Nevertheless, it is remarkable that $W^{\text{na}}$ describing all the non-adiabatic effects is such a simple and explicit functional of $\rho^{\rm S}$.
1D soft-Coulomb systems.
We next consider a two-electron singlet in 1D, with a soft-Coulomb interaction $V_{\text{ee}}(x)=1/\sqrt{x^{2}+1}$. This framework was extensively studied [33, 34, 35, 28, 36, 37], because standard adiabatic TDDFT approximations fail to reproduce, even qualitatively, the exact time-dependent density $\rho^{\rm S}$ for this 1D toy model in much the same way as they do for real 3D Coulombic systems. We consider two different time-dependent external potentials $V_{\rm ext}^{(1)}$ and $V_{\rm ext}^{(2)}$.
First, as in [33, 37], we study Rabi oscillations in a Helium-like atom perturbed by an electric dipole interaction, corresponding to
$V_{\rm ext}^{(1)}(t,x)=-\frac{2}{\sqrt{x^{2}+1}}+x\,E\,\sin(\omega t)$
where $E=0.00667$ and $\omega=0.5336$ is resonant with the first singlet excitation. The Rabi period is $T_{\rm R}=2\pi/(|d_{\rm eg}|E)$, where $d_{\rm eg}=\langle\Psi_{\rm e}|\widehat{x}_{1}+\widehat{x}_{2}|\Psi_{\rm g}\rangle$ and $\Psi_{\rm g}$ and $\Psi_{\rm e}$ are the ground and first excited states of $V_{\rm ext}^{(1)}(0,x)$.
To find a good approximation to $\rho^{\rm S}$, we follow common practice and solve a two-level time-dependent Schrödinger equation in the basis $(\Psi_{\rm g},\Psi_{\rm e})$. Those are obtained by diagonalizing a finite-difference approximation of the two-electron Hamiltonian at time $t=0$. The non-adiabatic contribution $V^{\text{na}}$ to the time-dependent exact Kohn–Sham potential is computed from (20) with $W=0$. In 1D, we have the explicit formula [38, 39]
$V^{\text{na}}:=V^{\rm eKS}-V^{\rm eaKS}=-\frac{1}{2}\left(\frac{j_{\varphi}}{\rho^{\rm S}}\right)^{2}-\int^{x}\partial_{t}\left(\frac{j_{\varphi}}{\rho^{\rm S}}\right),$
where $j_{\varphi}=\rho^{\rm S}\nabla\theta$ is the current. This potential is compared to the geometric non-adiabatic term $W^{\rm na}$ in (21). Our code is written in Julia and is freely available at [40].
In Fig. 3, we plot $V^{\rm na}$ and $W^{\rm na}$ over some intervals of times. Like in [28], we observe the presence of non-adiabatic fast-evolving steps in $V^{\text{na}}$ associated with a strong spatially nonlocal dependence on the density. In contrast, $W^{\rm na}$ is much more localized. Note also the scale difference between $V^{\text{na}}$ and $W^{\rm na}$.
Figure 2: Rabi oscillations in a 1D Helium-like atom with soft-Coulomb potential subjected to a time-dependent uniform electric field with amplitude $E=0.00667$ and angular frequency $\omega=0.5336$. Heatmaps of the non-adiabatic contribution $V^{\text{na}}$ to the exact time-dependent Kohn–Sham potential (left) and of the non-adiabatic geometric correction term $W^{\text{na}}$ (right), at times $|t-T_{\rm R}/4|\leqslant 2T_{\rm opt}$ (bottom) and $|t-T_{\rm R}/2|\leqslant 2T_{\rm opt}$ (top) with $T_{\rm opt}=2\pi/\omega$ the optical period.
Figure 3: Charge transfer in 1D between a Helium-like atom and a model of a closed-shell atom with frozen electrons with soft-Coulomb potential obtained by a time-dependent uniform electric field with amplitude $E=0.00667$ and angular frequency $\omega=0.112$. Heatmaps of the non-adiabatic contribution $V^{\text{na}}$ to the exact time-dependent Kohn–Sham potential (left) and of the non-adiabatic geometric correction term $W^{\text{na}}$ (right), at times $|t-T_{\rm R}/8|\leqslant 2T_{\rm opt}$ (bottom) and $|t-T_{\rm R}/4|\leqslant 2T_{\rm opt}$ (top) with $T_{\rm opt}=2\pi/\omega$ the optical period.
Next, we consider a charge-transfer mechanism from a Helium-type donor to a model of a closed-shell atom with frozen electrons, as in [36]:
$V_{\rm ext}^{(2)}(t,x)$ $=$ $-\frac{2}{\sqrt{\left(x+3.5\right)^{2}+1}}-\frac{1}{\cosh^{2}(x-3.5)}$
$+x\,E\,\sin(\omega t)$
with $E=0.00667$ and $\omega=0.112$. In Fig. 3, we plot $V^{\text{na}}$ and $W^{\text{na}}$ over intervals of times, computed numerically using the same approach as for the previous example. We find the same fast steps and high peaks in $V^{\rm na}$ as in [36]. In contrast, $W^{\text{na}}$ is more localized with a small amplitude and without high peaks.
Conclusion.
We have proposed a new TDDFT scheme based on geometric considerations. Our numerical simulations for the two-electron singlet suggest that approximating the non-adiabatic correction $W^{\rm na}$ could be easier than the corresponding non-adiabatic part of the standard Kohn–Sham potential, because it is a much smoother function of both time and space. Our new scheme could thus be more adapted to the description of out-of-equilibrium processes. In the companion article [10] we study the mathematical foundations of our geometric approach for finite-dimensional systems. Future work will be devoted to constructing practical approximations of the geometric correction $W$ for real systems.
Acknowledgments.
This work has benefited from French State support managed by ANR under the France 2030 program through the MaQuI CNRS Risky and High-Impact Research programme (RI)2 (grant agreement ANR-24-RRII-0001). We thank the referees for useful comments.
Appendix A Geometric Runge–Gross theorem
The Runge–Gross theorem [1] asserts that two time-dependent external potentials giving rise to the same density must necessarily be equal, up to a time-dependent additive constant. This is a fundamental result in TDDFT. Thus far, the mathematical proof of this result has only been possible under the assumption that the time-dependent potentials are spatially smooth and analytic in time and that the initial wavefunction is smooth [41]. The Runge–Gross argument therefore only applies to very smooth potentials such as the soft-Coulomb potential. Here we provide a rigorous proof that the same arguments as in [41] apply to our geometric theory. More precisely, we work in space dimension $d\in\{1,2,3\}$ and, to cover all situations of interest, we give ourselves four functions $V_{1}(t,\mathbf{r})$, $V_{2}(t,\mathbf{r})$, $W_{1}(t,\mathbf{r})$ and $W_{2}(t,\mathbf{r})$. We will suppose that either $V_{1}=V_{2}$ or $W_{1}=W_{2}$ and want to deduce the missing equality. We will always assume that
  • (i)
    the interaction $\widehat{V}_{\rm ee}$ is infinitely differentiable in space and all its derivatives are bounded;
  • (ii)
    the functions $V_{1}(t,\mathbf{r})$, $V_{2}(t,\mathbf{r})$, $W_{1}(t,\mathbf{r})$ and $W_{2}(t,\mathbf{r})$ are infinitely differentiable in space-time, with bounded space derivatives for all $t$;
  • (iii)
    the initial state $\Psi_{0}$ is infinitely differentiable in space and all its derivatives are square-integrable.
We also need a further assumption that will be different depending on the situation. We will require that the initial state $\Psi_{0}$ satisfies one of the two conditions:
  • (iv)
    $\rho_{\Psi_{0}}(\mathbf{r})>0$ everywhere;
  • (iv’)
    the unique $V$-representability property
    $\sum_{j=1}^{N}V(\mathbf{r}_{j})\Psi_{0}(\mathbf{r}_{1},...,\mathbf{r}_{N})=0\quad\Longrightarrow\quad V(\mathbf{r})=0$ (22)
    for all $V$ such that $\int_{\mathbb{R}^{d}}V(\mathbf{r})^{2}\rho_{\Psi_{0}}(\mathbf{r})\,\mathrm{d}\mathbf{r}<\infty$.
The equalities in (22) are understood Lebesgue-almost everywhere. The assumptions (i)–(iv) are from [41]. The unique $V$-representability property (22) in (iv’) is slightly stronger than (iv) and is further studied in [42] and in [10, Sec. VII.B]. This is the main property used in the proof of the Hohenberg–Kohn theorem for ground-state DFT, where it follows from the unique continuation principle [43, 44]. In particular, (22) is valid if $\Psi_{0}$ is the ground state of a Hamiltonian of the above form. For the theorem we do not need that $\Psi_{0}$ is a ground state but we do assume the condition (22). As explained in [10], this is the main property necessary for the geometric interpretation of TDDFT.
Theorem 1 (Geometric Runge–Gross).
Let $\Psi_{1}(t)$ and $\Psi_{2}(t)$ be the solutions to the Schrödinger equations
$i\partial_{t}\Psi_{m}(t)=\bigg(\widehat{T}+\widehat{V}_{\rm ee}+\sum_{j=1}^{N}V_{m}(t,\mathbf{r}_{j})+iW_{m}(t,\mathbf{r}_{j})\bigg)\Psi_{m}(t)$
with $m=1$ or $m=2$, starting at the same initial condition $\Psi_{0}$. We assume (i)–(iii) and that $\rho_{\Psi_{1}(t)}=\rho_{\Psi_{2}(t)}$ for $t\in[0,T)$.
  • (a)
    Assume that $W_{1}=W_{2}$, that (iv) holds and that the function $t\mapsto V_{2}(t,\mathbf{r})-V_{1}(t,\mathbf{r})$ is real-analytic for any $\mathbf{r}\in\mathbb{R}^{d}$. Then $V_{1}(t,\mathbf{r})=V_{2}(t,\mathbf{r})+C(t)$ for all $\mathbf{r}\in\mathbb{R}^{d}$ and all $t\in[0,T)$.
  • (b)
    Assume that $V_{1}=V_{2}$, that (iv’) holds and that the function $t\mapsto W_{2}(t,\mathbf{r})-W_{1}(t,\mathbf{r})$ is real-analytic for any $\mathbf{r}\in\mathbb{R}^{d}$. Then $W_{1}(t,\mathbf{r})=W_{2}(t,\mathbf{r})$ for all $\mathbf{r}\in\mathbb{R}^{d}$ and all $t\in[0,T)$.
We implicitly assume in the statement that $\Psi_{m}(t)$ stays normalized for all $t\in[0,T)$, that is, $\int_{\mathbb{R}^{d}}W_{m}(t,\mathbf{r})\rho_{\Psi_{m}(t)}(\mathbf{r})\,\mathrm{d}\mathbf{r}=0$. If we use the commutator form as in (9) then this implicit assumption is not necessary and uniqueness for $W$ holds only up to a time-dependent constant $C(t)$ like for $V$.
The proof of Theorem 1 follows closely [41] and we only explain it in case (b). Under the conditions (i)–(iii) and (iv’), we show that
$\frac{\partial^{\ell}}{\partial t^{\ell}}W_{1}(0,\mathbf{r})=\frac{\partial^{\ell}}{\partial t^{\ell}}W_{2}(0,\mathbf{r})$ (23)
for all $\ell$ and all $\mathbf{r}\in\mathbb{R}^{d}$. The conclusion then follows from the assumed analyticity in time. The smoothness of the potentials and $\Psi_{0}$ implies that we can freely differentiate Schrödinger’s equation in time [41]. We denote by $\widehat{W}_{m}(t):=\sum_{j=1}^{N}W_{m}(t,\mathbf{r}_{j})$ the many-body operators associated with $W_{1}$ and $W_{2}$. We also introduce the difference $W:=W_{2}-W_{1}$ which we want to prove vanishes, and the corresponding $\widehat{W}:=\widehat{W}_{2}-\widehat{W}_{1}$. We call $\rho(t,\mathbf{r})$ the common density of $\Psi_{1}(t)$ and $\Psi_{2}(t)$. Since in (b) we assume $V_{1}=V_{2}$, we call $H(t)=\widehat{T}+\widehat{V}_{\rm ee}+\widehat{V}_{1}(t)$ the common part of the Hamiltonian.
We then give ourselves a smooth time-independent function $U(\mathbf{r})$, that we will specify later. Using again the notation $\widehat{U}:=\sum_{j=1}^{N}U(\mathbf{r}_{j})$, we have
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{U}\Psi_{m}(t)\right\rangle}.$
Differentiating this relation once in time gives
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\partial_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}\\ =\Big\langle\Psi_{m}(t),\big(i[\widehat{T},\widehat{U}]+\widehat{U}\widehat{W}_{m}(t)+\widehat{W}_{m}(t)\widehat{U}\big)\Psi_{m}(t)\Big\rangle.$ (24)
We evaluate (24) at time $t=0$ and recall that $\Psi_{1}(0)=\Psi_{2}(0)=\Psi_{0}$. Subtracting the two equations (for $m=1$ and $m=2$) leads to
${\left\langle\Psi_{0},(\widehat{U}\widehat{W}(0)+\widehat{W}(0)\widehat{U})\Psi_{0}\right\rangle}=0$
where we recall that $W(0,\mathbf{r})=W_{2}(0,\mathbf{r})-W_{1}(0,\mathbf{r})$. It is natural to take $U(\mathbf{r})=W(0,\mathbf{r})$ so that the previous relation reduces to $\|\widehat{W}(0)\Psi_{0}\|^{2}=0$, that is, $\widehat{W}(0)\Psi_{0}=0$. From the unique $V$-representability property (22), this implies $W(0,\mathbf{r})=W_{2}(0,\mathbf{r})-W_{1}(0,\mathbf{r})=0$ almost everywhere, hence everywhere by continuity. This is the claimed result (23) for $\ell=0$.
Next we go on and differentiate once more (24), still with a general function $U$. To simplify the expression, we denote by $\widehat{L}_{m}^{(1)}(t):=i[\widehat{T},\widehat{U}]+\widehat{U}\widehat{W}_{m}(t)+\widehat{W}_{m}(t)\widehat{U}$ the Hermitian operator appearing in (24). The second-order derivative equals
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\partial^{2}_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{K}^{(1)}_{m}(t)\Psi_{m}(t)\right\rangle}\\ +{\left\langle\Psi_{m}(t),\big(\widehat{U}\partial_{t}\widehat{W}_{m}(t)+\partial_{t}\widehat{W}_{m}(t)\widehat{U}\big)\Psi_{m}(t)\right\rangle}$ (25)
with the operator
$\widehat{K}^{(1)}_{m}(t):=i[\widehat{H}(t),\widehat{L}^{(1)}_{m}(t)]+\widehat{L}_{m}^{(1)}(t)\widehat{W}_{m}(t)+\widehat{W}_{m}(t)\widehat{L}^{(1)}_{m}(t).$
We again evaluate at time $t=0$ and use that $\widehat{L}^{(1)}_{1}(0)=\widehat{L}^{(1)}_{2}(0)$ and $\widehat{K}^{(1)}_{1}(0)=\widehat{K}^{(1)}_{2}(0)$ since $\widehat{W}_{1}(0)=\widehat{W}_{2}(0)$, as we have proved in the first step. We find
${\left\langle\Psi_{0},\big(\widehat{U}\partial_{t}\widehat{W}(0)+\partial_{t}\widehat{W}(0)\widehat{U}\big)\Psi_{0}\right\rangle}=0.$
Taking $U(\mathbf{r})=\partial_{t}W(0,\mathbf{r})=\partial_{t}W_{2}(0,\mathbf{r})-\partial_{t}W_{1}(0,\mathbf{r})$ and using again (22) for that $U$, we arrive at the conclusion that $\partial_{t}W_{2}(0,\mathbf{r})=\partial_{t}W_{1}(0,\mathbf{r})$, which is (23) for $\ell=1$.
The proof goes on like this by induction. We have
$\int_{\mathbb{R}^{d}}U(\mathbf{r})\partial^{\ell}_{t}\rho(t,\mathbf{r})\,\mathrm{d}\mathbf{r}={\left\langle\Psi_{m}(t),\widehat{K}^{(\ell-1)}_{m}(t)\Psi_{m}(t)\right\rangle}\\ +{\left\langle\Psi_{m}(t),\big(\widehat{U}\partial^{\ell}_{t}\widehat{W}_{m}(t)+\partial^{\ell}_{t}\widehat{W}_{m}(t)\widehat{U}\big)\Psi_{m}(t)\right\rangle}$ (26)
for some operator $\widehat{K}^{(\ell-1)}_{m}$ defined by induction as
$\widehat{K}^{(\ell-1)}_{m}(t):=\partial_{t}\widehat{K}^{(\ell-2)}_{m}(t)+i\big[\widehat{H}(t),\widehat{L}^{(\ell-1)}_{m}(t)\big]\\ +\widehat{W}_{m}(t)\widehat{L}^{(\ell-1)}_{m}(t)+\widehat{L}^{(\ell-1)}_{m}(t)\widehat{W}_{m}(t),$
where
$\widehat{L}^{(\ell-1)}_{m}(t):=\widehat{K}^{(\ell-1)}_{m}(t)+\widehat{U}\partial^{\ell-1}_{t}\widehat{W}_{m}(t)+\partial^{\ell-1}_{t}\widehat{W}_{m}(t)\widehat{U}$.
The important remark is that $\widehat{K}_{m}^{(\ell-1)}(0)$ and $\widehat{L}_{m}^{(\ell-1)}(0)$ only depend on $\widehat{W}_{m}(0)$ and the derivatives $\partial^{k}_{t}\widehat{W}_{m}(0)$ up to order $k=\ell-1$, which coincide by the induction hypothesis.